Given the following information,

Real risk-free rate = 0.025

Inflation risk premium = 0.015

Maturity risk premium = 0.05

Default risk premium = 0.035

Liquidity risk premium = 0.01

(1) Using approximation method, what is the real rate of interest?

(2) Using Fisher equation, what is the real rate of interest?
Answer #1

Solution :-

Real Rate of interest Using Approximation method =

Real risk-free rate + Inflation risk premium + Maturity risk premium + Default risk premium + Liquidity risk premium = 0.025 + 0.015 + .05 + 0.035 + 0.01 = 0.09 = 9%

Real Rate of interest Using Fisher Equation =

(1+ r) = ( 1 + Real risk-free rate) * (1+ Inflation risk premium) * (1 + Maturity risk premium) * (1 + Default risk premium) * (1 + Liquidity risk premium ) =(1 + 0.025 ) * (1 + 0.015) * (1 + .05) * (1 + 0.035) * (1 + 0.01)

= (1+r) = (1.025)*(1.015)*(1.035)*(1.01) = 1.1419

therefore r = 0.1419 = 14.19%

Feel free to ask any query through comments as your satisfaction is very valuable.